OpenStudy (anonymous):

which equation represents a line that is parallel to the line y=3-2x and how? 1) 4x+2x=5 2) 2x+4y=1 3) y=3-4x 4)y=4x-2

OpenStudy (anonymous):

none of 'em...

OpenStudy (anonymous):

assuming you wrote those choices down correctly...

OpenStudy (anonymous):

my answer key says the answer is choice 1 but i have no idea how to get that.

OpenStudy (anonymous):

I think u need to write 1st option as 4x+2y=5 which is answer as it can be written as 2x+y=5/2 and hence parralel to y+2x=3

OpenStudy (anonymous):

ohhh yeah i wrote it wrong SO sorry! but how do i get that answer, meaning what steps do i take to figure that out?

OpenStudy (anonymous):

write everything in slope intercept form... if their slopes are the same they're parallel.

OpenStudy (anonymous):

as ax+by=c and ax+by=d are parralel and thats what i got in previous comment

OpenStudy (anonymous):

i hpe u get it now????

OpenStudy (anonymous):

not really...still a little confused!!

OpenStudy (anonymous):

in the form y= m*x+d m gives you the slope, meaning "if x increases by 1 , y must increase by m", if the m is equal for 2 or more lines, these are parallel, if d is also equal, then it is the same line

OpenStudy (anonymous):

Try to draw each one of them on a piece of squared paper, then try to understand the relation between the slope of a line and the coefficients of x and y. Also try to understand what the constant geometrically means.